Lionblaze18:

If one cup of flour can make 6 pancakes, how many pancakes can 7 cups of flour make?

Rylee88:

Well, do 6*7 and you will get your answer

Rylee88:

Because you're offline I will just explain it, so 6*7=42. 6*1=6 6*2=12 6*3=18 6*4=24 6*5=30 6*6=36 6*7=42 Or if you're not good at multiplication you can do this: 6+6+6+6+6+6+6=42. Hope this helps c:

IamIsis:

Hi so if 1 is to 6 then 2 is to 12 so you would multiply. 1x6=6 2x6=12 and keep going till you get to 7. Check 6x7=42

IamIsis: